# Large Language Model Supply Chain: A Research Agenda

### v2

@@ 4.2. Infrastructure: Datasets
The datasets employed for training play a crucial role in the development of LLMs (Jain et al., 2020; Sambasivan et al., 2021), which encompass vast collections of text from various sources, including books, websites, academic papers, social media, and open-source repositories. These datasets are complemented by data processing pipelines, annotation tools, and data management systems that enable efficient curation, cleaning, and utilization of the data. This vision aims to establish a data infrastructure that is diverse, scalable, inherently secure, and privacy-preserving (Mazumder et al., 2024).
@@
Current State. The current landscape of datasets in the LLM supply chain is characterized by unprecedented scale and diversity, coupled with growing concerns about data quality, bias, and privacy. Recent studies have shown that LLMs can perpetuate and amplify societal biases present in the source data (Kirk et al., 2021; Yu et al., 2024b; Navigli et al., 2023). Additionally, the inclusion of copyrighted material (Meeus et al., 2024; Li et al., 2024a) and privacy information (Carlini et al., 2021, 2023) in training datasets has also become a contentious legal issue. In the realm of Code LLMs, publicly available code repositories have become crucial sources of training data, presenting unique challenges related to code quality (Jain et al., 2023), licensing (Majdinasab et al., 2024; Katzy et al., 2024a), and potential security vulnerabilities (Yang et al., 2024) in the training data itself. What’s more, security vulnerabilities in data processing pipelines have emerged as a critical concern (Zhao et al., 2024c), highlighting the need for more rigorous security practices in dataset preparation and model training workflows. For instance, the discovery of CVE-2024-39705 in NLTK (Smartkeyss, 2024) and NVDB-CNVDB-2023879241 in HuggingFace (Alien, and Nicky, 2023) revealed potential arbitrary code execution vulnerabilities, posing significant risks in dataset processing. As the field continues to evolve rapidly, addressing these challenges in dataset management, security, and privacy remains crucial for the responsible development of future LLM technologies, both for natural language and code-related applications.
@@
Challenge I [SE + S&P Perspective]: Dataset Cleaning and Curation. The process of data cleaning and curation is a critical step in the development of LLMs, serving as the backbone for ensuring integrity, privacy, and ethical alignment. The primary obstacles stem from the handling of privacy (Carlini et al., 2021; Kim et al., 2024; Lukas et al., 2023), biased (Kirk et al., 2021; Yu et al., 2024b; Ferrara, 2023; Zhuo et al., 2023), and toxic (Weidinger et al., 2022; Zhuo et al., 2023; Ousidhoum et al., 2021) data in training sets, each of which presents unique challenge. The potential privacy challenges are twofold: ensuring that personally identifiable information (PII) is not present in the training data (Brown et al., 2022) and preventing the model from learning to reproduce or infer it from the patterns it is trained on (Carlini et al., 2021, 2023; Kim et al., 2024). Bias in training data is a well-documented issue that can lead models to perpetuate or even amplify existing prejudices (Steed et al., 2022; Bommasani et al., 2021; Liang et al., 2021). The challenge lies in identifying and mitigating biases, which are often deeply embedded in the data and reflective of broader societal biases (Gallegos et al., 2023; Steed et al., 2022). The presence of toxic and harmful content in training datasets poses a significant risk to the safety and reliability of LLMs (Weidinger et al., 2022; Zhuo et al., 2023; Ousidhoum et al., 2021; Birhane et al., 2024). Models trained on datasets containing such content may reproduce or even generate harmful outputs, undermining their applicability in diverse contexts. These challenges in data cleaning and curation require sophisticated strategies for mitigation, and this provides some opportunities as discussed below.
@@
  • Opportunity: Privacy Preserving. In the realm of privacy preservation for training LLMs, several techniques (e.g., text anonymization (Pilán et al., 2022) and differential privacy techniques (Hassan et al., 2019)) have emerged. However, challenges persist, particularly regarding knowledge integrity; removing sensitive information may lead to the loss of critical context, potentially misleading LLMs. Research focused on balancing anonymization and utility is essential. Additionally, the complexity of multimodal data complicates effective anonymization across diverse data types, highlighting the need for dedicated multimodal privacy frameworks.
@@
  • Opportunity: Detoxifying. Cleaning datasets of toxic content such as hate speech, misinformation, and abusive language requires not only sophisticated detection tools (Köpf et al., 2024; Zhang et al., 2024c; Bespalov et al., 2023) but also a nuanced understanding of the mechanism what constitutes harm (Welbl et al., 2021; Pavlopoulos et al., 2020; Baldini et al., 2021), which can vary widely across different cultural and social contexts. Cross-cultural sensitivity presents an opportunity to create guidelines and frameworks that respect cultural differences while identifying universally harmful content.
@@
  • Opportunity: Data Validation. The first line of defense against data poisoning is data validation (Polyzotis et al., 2019; Hutchinson et al., 2021), a process that is inherently complex due to the vast scale and heterogeneity of datasets used in LLM training. Effective validation requires sophisticated algorithms capable of detecting anomalies and malicious modifications in the data (Lwakatare et al., 2021), which is a task that becomes exponentially difficult as the data volume and diversity increase (Paleyes et al., 2022; Shivashankar and Martini, 2022). The opportunity for progress in robust data validation resides in advancing algorithmic solutions that are capable of nuanced detection of subtle and sophisticated data manipulation attempts. These solutions must be scalable enough to manage the expansive datasets characteristic of LLM training, thereby ensuring comprehensive coverage without compromising efficiency.
@@
  • Opportunity: Understanding Complex Licenses. One of the primary challenges in dataset licensing management is the complexity and diversity of licenses (Vendome and Poshyvanyk, 2016; Vendome et al., 2015, 2017a; Di Penta et al., 2010; Vendome, 2015). Data sources can range from openly accessible datasets with permissive licenses to proprietary datasets with stringent usage restrictions. Each source comes with its own set of legal terms, necessitating careful review and comprehension to ensure compliance (Wolter et al., 2023; Wu et al., 2017; BARCOMB and RIEHLE, 2022; German et al., 2010; Vendome et al., 2017b). Opportunities in this area include developing automated tools for detecting and summarizing key legal terms, providing stakeholders with clear and accessible insights into the permissions, obligations, and restrictions associated with each dataset.
@@ 4.3. Infrastructure: LLM Development Toolchain
Current State. The development toolchain for LLMs is characterized by a rapidly evolving ecosystem of third-party libraries, frameworks, and specialized tools designed to address the unique challenges of LLM development. As discussed in § 4.1 and § 4.2, distributed training frameworks and data processing pipelines introduce a set of security concerns. Beyond these, Hugging Face’s Transformers library has emerged as a de facto standard, providing a comprehensive suite of pre-trained models and fine-tuning tools. However, its widespread adoption has raised concerns about over-reliance on a single framework, potentially introducing systemic vulnerabilities across the LLM supply chain (Alien, and Nicky, 2023; Zhao et al., 2024c). Major deep learning frameworks like PyTorch and TensorFlow continue to play crucial roles, but they are not immune to security issues. For instance, TensorFlow’s Keras framework suffered from a Lambda Layer vulnerability (CVE-2024-3660) allowing arbitrary code injection (CERT Vulnerability Notes Database, 2024; Avi Lumelsky, 2024; Tom Bonner, 2023), while PyTorch’s use of Pickle for model serialization introduces potential deserialization vulnerabilities (Zhao et al., 2024c; Boyan Milanov, 2024a, b). The toolchain landscape has also seen the emergence of specialized tools addressing specific aspects of LLM development, such as Weights & Biases (Weights & Biases, 2024) for experiment tracking and Ray (Ray, 2024) for distributed computing. These tools, while enhancing productivity, introduce new potential attack vectors (Record, 2024; Shachar Menashe, 2024; Richard Harang, 2024). A recent path traversal vulnerability (CVE-2023-1177) in MLflow (iumiro, 2023), a popular experiment tracking tool, allowed unauthorized access to files outside the intended directory. This flaw risked exposing sensitive data and configuration files, underscoring the importance of proper input validation in LLM infrastructure tools. The increasing complexity of the toolchain, coupled with the rapid pace of development, has created challenges in maintaining consistent security standards across the LLM development pipeline.
@@
  • Opportunity: LLM Development Toolchain Mining. A promising avenue for enhancing supply chain security lies in the opportunity of LLM development toolchain mining. This approach involves the systematic analysis and evaluation (Jiang et al., 2022) of the tools and libraries used in the creation and training of LLMs. The core of this opportunity revolves around the comprehensive mining and auditing of development tools, from code libraries to data processing frameworks used in LLM training. Through the detailed analysis of the toolchain, developers can identify redundancies, inefficiencies, and areas for improvement, paving the way for the development of more streamlined, effective, and secure LLMs. Additionally, this mining process can spur innovation by highlighting gaps or needs within the toolchain, driving the creation of new tools, or the enhancement of existing ones to better serve the evolving demands of LLM development.
Challenge II [S&P Perspective]: Potential Toolchain Vulnerability. As the LLM development toolchain grows in complexity, conducting systematic vulnerability analyses becomes increasingly challenging (Wu et al., 2024b; Liu et al., 2023b; He et al., 2024b). Each component of the toolchain, from data preprocessing libraries to model serving platforms, introduces potential security risks. The challenge lies in developing comprehensive methodologies for identifying, assessing, and mitigating vulnerabilities across this diverse ecosystem. This includes addressing issues such as code injection vulnerabilities in AI frameworks, path traversal flaws in experiment tracking tools, and potential exploits in distributed training systems.
@@
Challenge III [S&P + SE Perspective]: Intricate Dependency and Vulnerability Propagation. Managing the intricate dependency tree, including both open-source and commercial components, poses a significant challenge within the supply chain (Liu et al., 2022a; Hejderup et al., 2018; Ma, 2018). Supply chain attacks targeting the development infrastructure or code repositories could lead to the injection of vulnerabilities or malicious code (Li et al., 2023d; Duan et al., 2020; Guo et al., 2023b), potentially compromising the entire lifecycle of model development and deployment. Moreover, vulnerabilities within dependencies or components can propagate through the supply chain (Hu et al., 2023; Zhang et al., 2023c), adversely affecting the security and reliability of models. Establishing robust dependency management processes, conducting thorough security monitoring, and ensuring supply chain transparency are essential for mitigating risks such as compromises in the LLM development tools supply chain.
  • Opportunity: Development of SBOM for LLM Toolchain. The adoption of the Software Bill of Materials (SBOM) of LLM toolchain presents a unique opportunity to achieve unprecedented levels of transparency and security. By meticulously documenting every library, dependency, and third-party component, SBOM enables developers to gain a comprehensive overview of their tools’ software ecosystem (Xia et al., 2023; Stalnaker et al., 2024; Stalnaker, 2023). This holistic visibility is instrumental in identifying vulnerabilities, outdated components, and non-compliant software elements that could jeopardize the development process and, ultimately, the security of the LLMs themselves. The detailed insights provided by SBOMs pave the way for proactive vulnerability management. Armed with knowledge about constituent components, development teams can swiftly address security flaws, apply necessary patches, and update components. This preemptive identification and remediation process is crucial in safeguarding models against potential exploits that could jeopardize their reliability and the security of the systems they operate within.
@@ 5.1. Model: Development & Training
Challenge [S&P + SE Perspective]: Inner and Outer Alignment. As the capabilities of LLMs continue to expand, ensuring their alignment (Liu et al., 2023g; Shen et al., 2023; Wolf et al., 2023) with human values and intentions becomes increasingly critical. The concept of inner alignment (Hubinger et al., 2019; Shen et al., 2023) focuses on ensuring that an LLM’s objectives are congruent with the intentions of its designers during the development and training phase. Outer alignment (Gossman and Kannan, 2023)—ensuring that the model’s outputs and behaviors align with societal norms and expectations—also plays a vital role in the overall effectiveness of LLMs. The pursuit of both inner and outer alignment requires a multifaceted strategy. Inner alignment is complicated by nuanced failure modes (Shen et al., 2023; Angelou, 2022; Hubinger et al., 2019), such as proxy, approximate, and suboptimality alignments, each presenting unique challenges in ensuring LLM systems operate as intended. These failure modes underscore the potential divergence between a model’s optimized objectives and the overarching goals its designers aim to achieve. Simultaneously, outer alignment challenges arise from the diverse values and expectations present in society, necessitating ongoing engagement with stakeholders to define acceptable behaviors and outputs. Methodological approaches such as relaxed adversarial training (Hubinger et al., 2019) and partitioning gradients (Yu et al., 2023) have been proposed to enhance inner alignment. However, achieving effective outer alignment requires transparency in the LLM system’s decision-making processes and continuous feedback from users and society at large.
@@
  • Opportunity: Advancing Interpretability of LLMs. The opportunity to advance the methodology of transparency and interpretability in LLMs stands as a critical endeavor (Singh et al., 2024). Enhancing transparency involves shedding light on the often opaque decision-making processes of these models, enabling a clearer understanding of how inputs are processed and interpreted to produce outputs (Zou et al., 2023a). By demystifying the inner workings of LLMs, researchers, and practitioners can gain valuable insights into the operational dynamics of these models, identifying areas where the models’ behaviors may not align with expected or desired outcomes (Singh et al., 2024; Jiang et al., 2021; He et al., 2024a). This early detection is invaluable, as it allows for timely interventions to correct course (Li et al., 2024e; Tan et al., 2024), preventing minor misalignments from escalating into more significant issues.
@@
  • Opportunity: Enhancing Feedback Mechanisms. The integration of robust feedback mechanisms into LLMs represents a transformative opportunity to enhance their adaptability and alignment with human values over time (Peng et al., 2023; Madaan et al., 2024). By embedding iterative feedback loops within the architecture of LLMs, developers can establish a dynamic process where the models continually learn and adjust from real-world interactions and user feedback. Feedback loops can be particularly beneficial in identifying and correcting biases, misconceptions, or inaccuracies that may emerge in LLM outputs, thereby enhancing the models’ trustworthiness and reliability (Ouyang et al., 2022; Peng et al., 2023; Liu et al., 2023f). This process enables LLMs to evolve and adapt in response to changing contexts, user needs, and societal norms, ensuring their ongoing relevance and utility.

### v3

@@ 4.2. Infrastructure: Datasets
The datasets employed for training play a crucial role in the development of LLMs (Jain et al., 2020; Sambasivan et al., 2021), which encompass vast collections of text from various sources, including books, websites, academic papers, social media, and open-source repositories. These datasets are complemented by data processing pipelines, annotation tools, and data management systems that enable efficient curation, cleaning, and utilization of the data. Our vision aims to establish a data infrastructure that is diverse, scalable, inherently secure, and privacy-preserving (Mazumder et al., 2023).
@@
Current State. The current landscape of datasets in the LLM supply chain is characterized by unprecedented scale and diversity, coupled with growing concerns about data quality, bias, and privacy. Recent studies have shown that LLMs can perpetuate and amplify societal biases present in the source data (Kirk et al., 2021; Yu et al., 2023b; Navigli et al., 2023). Additionally, the inclusion of copyrighted material (Meeus et al., 2024; Li et al., 2024a) and privacy information (Carlini et al., 2021, 2023) in training datasets has also become a contentious legal issue. In the realm of Code LLMs, publicly available code repositories have become crucial sources of training data, presenting unique challenges related to code quality (Jain et al., 2024b), licensing (Majdinasab et al., 2024; Katzy et al., 2024a), and potential security vulnerabilities (Yang et al., 2024) in the training data itself. What’s more, security vulnerabilities in data processing pipelines have emerged as a critical concern (Zhao et al., 2024b), highlighting the need for more rigorous security practices in dataset preparation and model training workflows. For instance, the discovery of CVE-2024-39705 in NLTK (Smartkeyss, 2024) and NVDB-CNVDB-2023879241 in Hugging Face (Alien, and Nicky, 2023) revealed potential arbitrary code execution vulnerabilities, posing significant risks in dataset processing. As the field continues to evolve rapidly, addressing these challenges in dataset management, security, and privacy remains crucial for the responsible development of future LLM technologies, both for natural language and code-related applications.
@@
Challenge I [SE + S&P Perspective]: Dataset Cleaning and Curation. The process of data cleaning and curation is a critical step in the development of LLMs, serving as the backbone for ensuring integrity, privacy, and ethical alignment. The primary obstacles stem from the handling of privacy (Carlini et al., 2021; Kim et al., 2023; Lukas et al., 2023), biased (Kirk et al., 2021; Yu et al., 2023b; Ferrara, 2023; Zhuo et al., 2023), and toxic (Weidinger et al., 2022; Zhuo et al., 2023; Ousidhoum et al., 2021) data in training sets, each of which presents unique challenge. The potential privacy challenges are twofold: ensuring that personally identifiable information (PII) is not present in the training data (Brown et al., 2022) and preventing the model from learning to reproduce or infer it from the patterns it is trained on (Carlini et al., 2021, 2023; Kim et al., 2023). Bias in training data is a well-documented issue that can lead models to perpetuate or even amplify existing prejudices (Steed et al., 2022; Bommasani et al., 2021; Liang et al., 2021). The challenge lies in identifying and mitigating biases, which are often deeply embedded in the data and reflective of broader societal biases (Gallegos et al., 2023; Steed et al., 2022). The presence of toxic and harmful content in training datasets poses a significant risk to the safety and reliability of LLMs (Weidinger et al., 2022; Zhuo et al., 2023; Ousidhoum et al., 2021; Birhane et al., 2023). Models trained on datasets containing such content may reproduce or even generate harmful outputs, undermining their applicability in diverse contexts. These challenges in data cleaning and curation require sophisticated strategies for mitigation, and this provides some opportunities as discussed below.
@@
  • Opportunity: Privacy Preserving. In the realm of privacy preservation for training LLMs, several techniques (e.g., text anonymization (Pilán et al., 2022) and differential privacy techniques (Hassan et al., 2020)) have emerged. However, challenges persist, particularly regarding knowledge integrity; removing sensitive information may lead to the loss of critical context, potentially misleading LLMs. Research focused on balancing anonymization and utility is essential. Additionally, the complexity of multimodal data complicates effective anonymization across diverse data types, highlighting the need for dedicated multimodal privacy frameworks.
@@
  • Opportunity: Detoxifying. Cleaning datasets of toxic content such as hate speech, misinformation, and abusive language requires not only sophisticated detection tools (Köpf et al., 2023; Zhang et al., 2024c; Bespalov et al., 2023) but also a nuanced understanding of the mechanism what constitutes harm (Welbl et al., 2021; Pavlopoulos et al., 2020; Baldini et al., 2022), which can vary widely across different cultural and social contexts. Cross-cultural sensitivity presents an opportunity to create guidelines and frameworks that respect cultural differences while identifying universally harmful content.
@@
  • Opportunity: Data Validation. The first line of defense against data poisoning is data validation (Breck et al., 2019; Hutchinson et al., 2021), a process that is inherently complex due to the vast scale and heterogeneity of datasets used in LLM training. Effective validation requires sophisticated algorithms capable of detecting anomalies and malicious modifications in the data (Lwakatare et al., 2021), which is a task that becomes exponentially difficult as the data volume and diversity increase (Paleyes et al., 2023; Shivashankar and Martini, 2022). The opportunity for progress in robust data validation resides in advancing algorithmic solutions that are capable of nuanced detection of subtle and sophisticated data manipulation attempts. These solutions must be scalable enough to manage the expansive datasets characteristic of LLM training, thereby ensuring comprehensive coverage without compromising efficiency.
@@
  • Opportunity: Understanding Complex Licenses. One of the primary challenges in dataset licensing management is the complexity and diversity of licenses (Vendome and Poshyvanyk, 2016; Vendome et al., 2015, 2017a; Penta et al., 2010; Vendome, 2015). Data sources can range from openly accessible datasets with permissive licenses to proprietary datasets with stringent usage restrictions. Each source comes with its own set of legal terms, necessitating careful review and comprehension to ensure compliance (Wolter et al., 2023a; Wu et al., 2017; Wolter et al., 2023b; Germán et al., 2010; Vendome et al., 2017b). Opportunities in this area include developing automated tools for detecting and summarizing key legal terms, providing stakeholders with clear and accessible insights into the permissions, obligations, and restrictions associated with each dataset.
@@ 4.3. Infrastructure: LLM Development Toolchain
Current State. The development toolchain for LLMs is characterized by a rapidly evolving ecosystem of third-party libraries, frameworks, and specialized tools designed to address the unique challenges of LLM development. As discussed in § 4.1 and § 4.2, distributed training frameworks and data processing pipelines introduce a set of security concerns. Beyond these, Hugging Face’s Transformers library has emerged as a de facto standard, providing a comprehensive suite of pre-trained models and fine-tuning tools. However, its widespread adoption has raised concerns about over-reliance on a single framework, potentially introducing systemic vulnerabilities across the LLM supply chain (Alien, and Nicky, 2023; Zhao et al., 2024b). Major deep learning frameworks like PyTorch and TensorFlow continue to play crucial roles, but they are not immune to security issues. For instance, TensorFlow’s Keras framework suffered from a Lambda Layer vulnerability (CVE-2024-3660) allowing arbitrary code injection (CERT Vulnerability Notes Database, 2024; Avi Lumelsky, 2024; Tom Bonner, 2023), while PyTorch’s use of Pickle for model serialization introduces potential deserialization vulnerabilities (Zhao et al., 2024b; Boyan Milanov, 2024a, b). The toolchain landscape has also seen the emergence of specialized tools addressing specific aspects of LLM development, such as Weights & Biases (Weights & Biases, 2024) for experiment tracking and Ray (Ray, 2024) for distributed computing. These tools, while enhancing productivity, introduce new potential attack vectors (Record, 2024; Shachar Menashe, 2024; Richard Harang, 2024). A recent path traversal vulnerability (CVE-2023-1177) in MLflow (iumiro, 2023), a popular experiment tracking tool, allowed unauthorized access to files outside the intended directory. This flaw risked exposing sensitive data and configuration files, underscoring the importance of proper input validation in LLM infrastructure tools. The increasing complexity of the toolchain, coupled with the rapid pace of development, has created challenges in maintaining consistent security standards across the LLM development pipeline.
@@
  • Opportunity: LLM Development Toolchain Mining. A promising avenue for enhancing supply chain security lies in the opportunity of LLM development toolchain mining, which involves the systematic analysis and evaluation (Jiang et al., 2022) of the tools and libraries used in the lifecycle of LLMs. The core of this opportunity revolves around the comprehensive mining and auditing of development tools, from code libraries to data processing frameworks used in LLM training. Through the detailed analysis of the toolchain, developers can identify redundancies, inefficiencies, and areas for improvement, paving the way for the development of more streamlined, effective, and secure LLMs. Additionally, this mining process can spur innovation by highlighting gaps or needs within the toolchain, driving the creation of new tools, or the enhancement of existing ones to better serve the evolving demands of LLM development.
Challenge II [S&P Perspective]: Potential Toolchain Vulnerability. As the LLM toolchain grows in complexity, conducting systematic vulnerability analyses becomes increasingly challenging (Wu et al., 2024b; Liu et al., 2023b; He et al., 2024b). Each component of the toolchain, from data preprocessing to model serving, introduces potential security risks. The challenge lies in developing comprehensive methodologies for identifying, assessing, and mitigating vulnerabilities across this diverse ecosystem.
@@
Challenge III [S&P + SE Perspective]: Intricate Dependency and Vulnerability Propagation. Managing the intricate dependency tree, including both open-source and commercial components, poses a significant challenge within the supply chain (Liu et al., 2022a; Hejderup et al., 2018; Ma, 2018). Supply chain attacks targeting the development infrastructure or code repositories could lead to the injection of vulnerabilities or malicious code (Li et al., 2023e; Duan et al., 2021; Guo et al., 2023b), potentially compromising the entire lifecycle of model development and deployment. Moreover, vulnerabilities within dependencies or components can propagate through the supply chain (Hu et al., 2024; Zhang et al., 2023c), adversely affecting the security and reliability of models. Establishing robust dependency management processes, conducting thorough security monitoring, and ensuring supply chain transparency are essential for mitigating risks such as compromises in the LLM development tools supply chain.
  • Opportunity: Development of SBOM for LLM Toolchain. The adoption of the software bill of materials (SBOM) of LLM toolchain presents a unique opportunity to achieve unprecedented levels of transparency and security. By meticulously documenting every library, dependency, and third-party component, SBOM enables developers to gain a comprehensive overview of their tools’ software ecosystem (Xia et al., 2023; Stalnaker et al., 2024). This holistic visibility is instrumental in identifying vulnerabilities, outdated components, and non-compliant software elements that could jeopardize the development process and, ultimately, the security of the LLMs themselves. The detailed insights provided by SBOMs pave the way for proactive vulnerability management. Armed with knowledge about constituent components, development teams can swiftly address security flaws, apply necessary patches, and update components. This preemptive identification and remediation process is crucial in safeguarding models against potential exploits that could jeopardize their reliability and the security of the systems they operate within.
@@ 5.1. Model: Development & Training
Challenge [S&P + SE Perspective]: Model Alignment. As the capabilities of LLMs continue to expand, ensuring their alignment (Liu et al., 2023g; Shen et al., 2023; Wolf et al., 2024) with human values and intentions become increasingly critical. The concept of inner alignment (Hubinger et al., 2019; Shen et al., 2023) focuses on ensuring that an LLM’s objectives are congruent with the intentions of its designers during the development and training phase. Outer alignment (Shen et al., 2023)—ensuring that the model’s outputs and behaviors align with societal norms and expectations—also plays a vital role in the overall effectiveness of LLMs. The pursuit of both inner and outer alignment requires a multifaceted strategy. Inner alignment is complicated by nuanced failure modes (Shen et al., 2023; Angelou, 2022; Hubinger et al., 2019), such as proxy, approximate, and suboptimality alignments, each presenting unique challenges in ensuring LLM systems operate as intended. These failure modes underscore the potential divergence between a model’s optimized objectives and the overarching goals its designers aim to achieve. Simultaneously, outer alignment challenges arise from the diverse values and expectations present in society, necessitating ongoing engagement with stakeholders to define acceptable behaviors and outputs. Methodological approaches such as relaxed adversarial training (Hubinger et al., 2019) and partitioning gradients (Yu et al., 2023a) have been proposed to enhance inner alignment. However, achieving effective outer alignment requires transparency in the LLM system’s decision-making processes and continuous feedback from users and society at large.
@@
  • Opportunity: Advancing Interpretability of LLMs. The opportunity to advance the methodology of transparency and interpretability in LLMs stands as a critical endeavor (Singh et al., 2024). Enhancing transparency involves shedding light on the often opaque decision-making processes of these models, enabling a clearer understanding of how inputs are processed and interpreted to produce outputs (Zou et al., 2023a). By demystifying the inner workings of LLMs, researchers, and practitioners can gain valuable insights into the operational dynamics of these models, identifying areas where the models’ behaviors may not align with expected or desired outcomes (Singh et al., 2024; Jiang et al., 2021; He et al., 2024a). This early detection is invaluable, as it allows for timely interventions to correct course (Li et al., 2023d; Tan et al., 2024), preventing minor misalignments from escalating into more significant issues.
@@
  • Opportunity: Enhancing Feedback Mechanisms. The integration of robust feedback mechanisms into LLMs represents a transformative opportunity to enhance their adaptability and alignment with human values over time (Peng et al., 2023; Madaan et al., 2023b). By embedding iterative feedback loops within the architecture of LLMs, developers can establish a dynamic process where the models continually learn and adjust from real-world interactions and user feedback. Feedback loops can be particularly beneficial in identifying and correcting biases, misconceptions, or inaccuracies that may emerge in LLM outputs, thereby enhancing the models’ trustworthiness and reliability (Ouyang et al., 2022; Peng et al., 2023; Liu et al., 2024b). This process enables LLMs to evolve and adapt in response to changing contexts, user needs, and societal norms, ensuring their ongoing relevance and utility.